OpenStudy (anonymous):

True or False This set of three numbers are measurements of a right triangle... 70,48,24

Parth (parthkohli):

false

OpenStudy (anonymous):

8,7,5 what about these???

Parth (parthkohli):

Yes. 49 + 25 = 64

OpenStudy (anonymous):

can you please explain why? so i understand

OpenStudy (anonymous):

how did you get 49 and 25?

Parth (parthkohli):

a^2 + b^2 = c^2(c is the largest side aka hypotenuse) => 7^2 + 5^2 = 8^2 => 49 + 25 = 64 => 64 = 64

OpenStudy (anonymous):

thank you so much

OpenStudy (zarkon):

if you like ParthKohli's reply you should give him/her a 'Best Answer'

Parth (parthkohli):

Thanks Zarkon :)